A tin which is 12cm long,9 cm wide and 4cm deep holds 120 g of tea. If 1 kg of the same tea is packed into a tin which has a 12 cm square base, how tall will the tin have to be?
Answer by rothauserc(4718)

volume of tin1 is 12 * 9 * 4 = 432 cubic cm which holds 120 g of tea
:
Note 1 Kg = 1000 g
:
120 / 432 = 1000 / x
:
120x = 432,000
:
x = 3600 cubic cm
:
volume of tin2 = 3600 = 12^2 * height
:
height = 3600 / 144 = 25
:
**********************
The tin2 is 25 cm tall
**********************
: